OpenStudy (anonymous):

a man cover 27kms in 9hours abd return home with a speed of 5km/hr.what is his average speed for entire journey

OpenStudy (mayankdevnani):

first of all. what is the speed at first situation, given that , d=27km and t=9hr and can you find the speed??

OpenStudy (mayankdevnani):

\[\huge speed=\frac{distance \space travelled}{time \space taken}\]

OpenStudy (mayankdevnani):

can you now solve it!! @qaiser380

OpenStudy (anonymous):

for average speed there is a simple relation av. speed = v = total distance travelled/ total time taken and hence the average speed would be 27/9 = 3 km/hr and in SI system 0.833... m/s

OpenStudy (anonymous):

He goes with a speed of 27/9 = 3km/hr. He returns with a speed of 5km/hr. So his average speed = (3+5)/2 = 4km/hr.

OpenStudy (mayankdevnani):

@rijusarkarr plzz don't give direct answers. let @qaiser380 be understand!